OpenStudy (anonymous):

In triangle ABC, what is the measure of angle C if A = 100°, a = 125, and c = 10? 4.52° 8° 12.31° Angle C does not exist; no triangle can be formed.

OpenStudy (dtan5457):

Use the laws of sin sinA/a=sinC/c

OpenStudy (dtan5457):

*sine

OpenStudy (dtan5457):

remember when your looking for an angle, you have to use sin^-1

OpenStudy (anonymous):

so how am I going to set the formula up?

OpenStudy (dtan5457):

\[\frac{ \sin100 }{ 125 }=\frac{ sinC }{ 10 }\]

OpenStudy (dtan5457):

use a calculator and solve for sin C

OpenStudy (anonymous):

so its 12.31

OpenStudy (dtan5457):

no you forgot to plug in sin^-1

OpenStudy (dtan5457):

sin^-1(sin100(10)/(125)=angle c

OpenStudy (anonymous):

yup I knew it 4.519

OpenStudy (dtan5457):

good job